A wave is modeled by the wave function y(x,t) = (0.51m) sin[(2π/7.9m)(x-(14.3m/s)t)]. What is the period of the wave?




Round your answer to 2 decimal places.



Pls dont forget the unit

1
Expert's answer
2022-05-15T17:03:24-0400
"y(x,t) =y_0\sin[(2π/\lambda)(x-vt)]""y(x,t) = (0.51m) \sin[(2π/7.9m)(x-(14.3m/s)t)]"

The period of wave



"T=\lambda/v=(7.9\:\rm m)/(14.3\:\rm m/s)=0.55\:\rm s"